### v1

# Conflicts in Texts: Data, Implications and Challenges

Abstract
As NLP models become increasingly integrated into real-world applications, it becomes clear that there is a need to address the fact that models often rely on and generate conflicting information. Conflicts could reflect the complexity of situations, changes that need to be explained and dealt with, difficulties in data annotation, and mistakes in generated outputs. In all cases, disregarding the conflicts in data could result in undesired behaviors of models and undermine NLP models’ reliability and trustworthiness. This survey categorizes these conflicts into three key areas: (1) natural texts on the web, where factual inconsistencies, subjective biases, and multiple perspectives introduce contradictions; (2) human-annotated data, where annotator disagreements, mistakes, and societal biases impact model training; and (3) model interactions, where hallucinations and knowledge conflicts emerge during deployment. While prior work has addressed some of these conflicts in isolation, we unify them under the broader concept of conflicting information, analyze their implications, and discuss mitigation strategies. We highlight key challenges and future directions for developing conflict-aware NLP systems that can reason over and reconcile conflicting information more effectively.
1 Introduction
The rapid advancement of natural language processing (NLP), particularly with the rise of large language models (LLMs), has led to their widespread adoption in daily tasks, information retrieval, and decision-making processes. However, the increasing complexity of these models reveals various types of conflicts at multiple stages, including training, annotation, and model interaction, affecting the reliability and trustworthiness of downstream applications. For example, training models on data containing factual contradictions, annotation disagreements, or prompts that contradict a model’s parametric knowledge can introduce inconsistencies with unpredictable consequences Pavlick and Kwiatkowski (2019); Sap et al. (2019).
Figure 1: Examples of the three different ares of conflicts discussed in this work. The first example describes a case where two different entities of the same name are found naturally on the web, the second example elaborates the annotation disagreement in a sentiment analysis task, and the third showcases a knowledge conflict between the context and memory of LLMs during model interactions.
{forest}
for tree= grow=east, parent anchor=east, child anchor=west, edge=draw, thick, rounded corners, draw=black, fit=band, minimum width=5cm, align=center, anchor=west, s sep=18pt, l sep=22pt, inner sep=10pt [Conflicts in NLP, fill=yellow!40, tier=0 [Model Interactions, fill=yellow!30, tier=1 [Hallucination, fill=green!40, tier=2 [Contextual Hallucination, fill=red!40, tier=3 [Maynez et al. (2020),
Raunak et al. (2021), Dale et al. (2023), fill=cyan!40] ] [Factual Hallucination, fill=red!40, tier=3 [Lin et al. (2022),
Pagnoni et al. (2021), Honovich et al. (2021), fill=cyan!40] ] ] [Knowledge Conflicts, fill=green!40, tier=2 [Across Different Models, fill=red!40, tier=3 [Cohen et al. (2023),
Zhu et al. (2024), Zhao et al. (2024), fill=cyan!40] ] [Between Context and Parameters, fill=red!40, tier=3 [Longpre et al. (2021); Chen et al. (2022a),
Chen et al. (2021a); Lazaridou et al. (2021), fill=cyan!40] ] ] ] [Human Annotations, fill=yellow!30, tier=1 [Ethical and Societal Biases, fill=green!40, tier=2 [Annotation Biases, fill=red!40, tier=3 [Sap et al. (2022), Faisal et al. (2022),
Thorn Jakobsen et al. (2022) , fill=cyan!40] ] ] [Human Judgements, fill=green!40, tier=2 [Interpretation Disagreements, fill=red!40, tier=3 [Kahneman (2021), Uma et al. (2021),
Sandri et al. (2023), Jiang and de Marneffe (2022), fill=cyan!40] ] ] ] [Natural Conflicts on Web, fill=yellow!30, tier=1 [Conflicts in Opinions, fill=green!40, tier=2 [Framing Bias, fill=red!40, tier=3 [Card et al. (2015), Liu et al. (2019),
Fan et al. (2019), Lei et al. (2022), fill=cyan!40] ] [Perspectives, fill=red!40, tier=3 [Chen et al. (2019), Liu et al. (2021a),
Lee et al. (2022) , Wan et al. (2024) , fill=cyan!40] ] ] [Factual Conflicts, fill=green!40, tier=2 [Contradictory Evidence, fill=red!40, tier=3 [Chen et al. (2022a),Hong et al. (2024),
Liu et al. (2024), Pham et al. (2024), fill=cyan!40] ] [Ambiguity, fill=red!40, tier=3 [Min et al. (2020), Zhang and Choi (2021),
Dhingra et al. (2022), Cole et al. (2023), fill=cyan!40] ] ] ] ]
Figure 2: Taxonomy of conflicts in texts.
Existing work on conflicts in NLP tends to focus on specific issues, such as annotation disagreements Uma et al. (2021), hallucinations and factuality Zhang et al. (2023); Wang et al. (2023), and knowledge conflicts Xu et al. (2024); Feng et al. (2024), without synthesizing these problems into a broader perspective. In this survey, we conceptualize these diverse challenges under the umbrella of conflicting information and analyze their origins, implications, and mitigation strategies. We first examine conflicts inherent in training data, including natural texts from the web and human-annotated datasets. We then explore conflicts arising during interactions with models in the LLM era, discussing their impact on downstream tasks and highlighting key challenges and future directions for conflict-aware AI systems.
The abundance of online data is accompanied by inherent conflicts, stemming from diverse sources, interpretations, and biases. These conflicts manifest as factual conflicts, such as semantic ambiguities Pavlick and Tetreault (2016); Min et al. (2020) and factual inconsistencies Pham et al. (2024); Liu et al. (2024), or as conflicts in opinions related to political ideologies Entman (1993); Recasens and et al. (2013) and perspectives Chen et al. (2019); Liu et al. (2021a). Factual conflicts are particularly prevalent in open-domain question answering (QA) and retrieval-augmented generation (RAG) systems Chen et al. (2017), where aggregating knowledge from multiple sources introduces inconsistencies Liu et al. (2024). These challenges highlight the need for conflict-aware retrieval and reasoning mechanisms to improve model reliability Xie et al. (2024). Unlike factual conflicts, opinionated disagreements reflect the variability in human interpretation, beliefs, and ideological stances Chen et al. (2019); Fan et al. (2019). The presence of conflicting viewpoints complicates tasks such as summarization, sentiment analysis, and dialogue generation, where maintaining coherence and neutrality is crucial Liu et al. (2021a); Lee et al. (2022). Furthermore, the uneven distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2024).
Another significant conflict arises in human-annotated data. For instance, annotation disagreements persists in both subjective and seemingly objective NLP tasks Mostafazadeh Davani et al. (2022). Disagreements are widespread in sentiment analysis Wan et al. (2023), hate speech detection Sap et al. (2022), and even natural language inference (NLI) Pavlick and Kwiatkowski (2019). Models trained on aggregated (e.g. majority-vote) labels struggle with ambiguous or high-disagreement examples, often treating them as hard-to-learn or mislabeled Anand et al. (2023). Pavlick and Kwiatkowski (2019) also find that standard NLI models’ uncertainty does not reflect the true ambiguity present in human opinions, leading to overconfidence in contentious cases. In addition, annotation biases—such as those related to race, gender, and geography—skew model predictions and reinforce societal biases Buolamwini and Gebru (2018); Sap et al. (2022); Pei and Jurgens (2023). These issues highlight the need for fair and representative annotations that capture the complexity of human disagreement.
Conflicts also emerge during interactions with models, manifesting as knowledge conflicts between model memories and contexts, and hallucinations in generated outputs. Knowledge conflicts arise when a model’s internal memory contradicts external contextual evidence, as shown by Longpre et al. (2021), who found that models often overly depend on memorized knowledge, leading to hallucinations. Neeman et al. (2023) proposed separating parametric and contextual knowledge to improve interpretability, while Xie et al. (2024) examined LLMs’ confirmation bias, showing how models inconsistently handle contradictory evidence. Additionally, hallucinations—ranging from factual inconsistencies Lin et al. (2022); Ouyang et al. (2022) to contextual hallucinations Maynez et al. (2020); Kryscinski et al. (2020)—further undermine model reliability. Various mitigation strategies have been proposed, including retrieval augmentation Lewis et al. (2020); Shuster et al. (2021), hallucination detection Manakul et al. (2023), and knowledge graph-based verification Guan et al. (2024).
By systematically categorizing and analyzing these conflicts, this survey provides a unified perspective on their origins, implications, and mitigation strategies. Addressing these issues is essential for building robust and trustworthy AI systems that operate effectively across diverse domains and user groups. Our findings contribute to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
2 Conflicts in Natural Texts on the Web
Conflicts in natural texts on the web manifest in diverse ways, reflecting the inherent complexity and subjectivity of human language. They can broadly be categorized into factual conflicts, which revolve around factual discrepancies caused by various reasons, and conflicts in opinions, which pertain to divergent perspectives or biases.
2.1 Factual Conflicts
2.1.1 Origins
Ambiguity
Ambiguity is a root cause of factual conflict. When a query or piece of data lacks clarity about entities or context, a model can produce conflicting answers. A clear demonstration of how ambiguity induces conflicts is context dependence. For example, an ambiguous question of "which COVID-19 vaccine was the first to be authorized by our government?" can have conflicting answers depending on different geographical contexts Zhang and Choi (2021).
Min et al. (2020) was the first work to study the effects of ambiguity in open domain question answering. They introduced AmbigQA, a dataset highlighting that over half of the open-domain, natural questions are ambiguous, with diverse sources of ambiguity such as event and entity references. Zhang and Choi (2021) proposed the SituatedQA task, showing that a significant fraction of open-domain questions are valid only under particular temporal or geographic contexts. Many other work specifically focus on the temporal aspect of ambiguity, benchmarking and evaluating models’ awareness and adaptation to time-sensitive questions Chen et al. (2021b); Liska et al. (2022); Kasai et al. (2023).
Contradictory Evidence
Conflicts in NLP systems arise when information on the web presents conflicting evidence towards a factual question. This issue is particularly prevalent in open-domain question answering settings, where models must navigate inconsistencies across diverse information sources. Liu et al. (2024) find that 25% of unambiguous factual questions queried on Google retrieve conflicting evidence from multiple sources. For instance, a Google search for "When was Kendrick Lamar’s first album released?" yields conflicting evidence, illustrating the challenge of integrating contradictory information in retrieval-based QA systems Liu et al. (2024).
Researchers have proposed different datasets to systematically study how NLP models handle such conflicts. Li et al. (2024b) introduce ContraDoc, a human-annotated dataset of long documents with internal contradictions; Pham et al. (2024) propose WhoQA, a benchmark dataset that constructs conflicts by formulating questions about a shared property among entities with the same name (e.g. "Who is George Washington?"); and Liu et al. (2024) construct QACC, a human-annotated dataset of conflicting results retrieved by Google. Beyond empirical datasets, several studies have proposed synthetic approaches to simulate conflicts through entity substitution Chen et al. (2022a); Hong et al. (2024), machine-generated conflicting evidence Pan et al. (2023); Wan et al. (2024); Hong et al. (2024), and pre-defined rule-based templates Kazemi et al. (2023).
2.1.2 Implications and Mitigation
Zhang and Choi (2021) show that pre-trained language models perform competitively at identifying whether a question is context dependent, matching human agreements, but lag behind human-level performance by a significant margin when answering questions dependent on temporal contexts. Cole et al. (2023) demonstrate that large language models benefit from a “disambiguate-then-answer” pipeline, showing improved reliability by detecting ambiguity before attempting an answer. Dhingra et al. (2022) propose time-aware language models that condition on timestamps to mitigate confusion arising from outdated facts or evolving knowledge.
Studies have shown that retrieval performance heavily impacts which sources models rely on Chen et al. (2022a), knowledge conflicts within contexts significantly degrade LLMs’ performance in RAG settings Pham et al. (2024); Liu et al. (2024); Li et al. (2024b), and QA models are vulnerable to even small amounts of evidence contamination brought by misinformation Pan et al. (2023). In addition, large language models (LLMs) exhibit a strong confirmation bias, favoring external evidence that aligns with their parametric memory, even when conflicting evidence is present Xie et al. (2024). To mitigate these issues, Hong et al. (2024) proposed fine-tuning a discriminator or prompting GPT-3.5 to elicit its discriminative capability and show that these approaches significantly enhance model robustness. On the other hand, Liu et al. (2024) proposed fine-tuning LLMs with human-written explanations to teach models to reason through conflicting evidence.
2.2 Conflicts in Opinions
2.2.1 Origins
Perspectives
Individuals and communities often hold diverse perspectives on the same issue. Such diversity is evident in online discussions and debates, where the multiplicity of viewpoints can lead to conflicting opinions. For instance, on controversial topics such as "Animals should have lawful rights," people express varying stances Chen et al. (2019), posing challenges for downstream tasks like summarization where consolidating viewpoints and presenting unbiased information are crucial Liu et al. (2021a); Lee et al. (2022).
Chen et al. (2019) address this challenge by presenting a range of perspectives on a given claim. The authors introduce the task of substantiated perspective discovery, where a system identifies diverse, evidence-supported perspectives that take a stance on a claim, and curated a dataset, PERSPECTRUM, for the task, using online debate platforms and search engines. Wan et al. (2024) introduce ConflictingQA, a dataset comprising controversial questions paired with real-world evidence documents presenting divergent facts, argumentation styles, and conclusions. Plepi et al. (2024) study perspective-taking in contentious online conversations (e.g. social media dilemmas). The authors create a new corpus of 95k conflict scenarios augmented with each user’s self-disclosed background information. Liu et al. (2021a) propose MultiOpEd, an open-domain corpus focusing on automatic perspective discovery. MultiOpEd comprises 1,397 controversial topics, each accompanied by two editorials expressing opposing viewpoints, and each editorial includes a one-sentence perspective summarizing its core argument and a brief abstract highlighting supporting details.
Framing Bias
A specific example of how differing opinions are conveyed and expanded is framing bias, a mechanism in which news media shape interpretations by emphasizing certain aspects of information over others Entman (1993). In a polarized media environment, partisan media outlets deliberately frame news stories in a way to advance certain political ideologies Jamieson et al. (2007); Levendusky (2013); Liu et al. (2019).
Various studies have examined different facets of such media bias. Card et al. (2015) introduce the Media Frames Corpus (MFC) to facilitate the computational study of media framing in news articles. The MFC comprises several thousand news articles covering three policy issues, each annotated according to 15 general-purpose framing dimensions. Liu et al. (2019) introduce Gun Violence Frame Corpus (GVFC), a dataset of news headlines with frames curated and annotated by journalism and communication experts. Fan et al. (2019) focus on informational bias, where factual content is subtly framed by content selection and organization within articles. They introduce BASIL, a dataset comprising 300 news articles annotated with 1,727 bias spans, revealing that informational bias occurs more frequently than lexical bias. Lei et al. (2022) identify ideological bias at the sentence level by analyzing news discourse structure, showing that bias-indicative sentences may appear neutral in isolation.
2.2.2 Implications and Mitigation
Analysis of PERSPECTRUM reveals significant natural language understanding challenges, as human performance substantially outperforms machine baselines at identifying diverse, evidence-supported perspectives Chen et al. (2019). Furthermore, when selecting real-world evidence for controversial questions, LLMs predominantly prioritize the relevance of the evidence to the query, often disregarding stylistic attributes such as the presence of scientific references or a neutral tone Wan et al. (2024). In addition, the distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al. (2023); Mihalcea et al. (2024). Studies have shown that LLMs’ outputs skew toward the values of Western English-speaking countries Tao et al. (2024); Naous et al. (2024), and misalignment is more pronounced for underrepresented personas and on culturally sensitive topics such as social values Al Kuwatly et al. (2020). Furthermore, LLMs often provide inconsistent answers to the same question when prompted in different languages Li et al. (2024a); AlKhamissi et al. (2024), revealing conflicting cultural perspectives within a single model.
Several studies have proposed methods to address conflicts in perspectives. Liu et al. (2021a) show that incorporating auxiliary tasks enhances the quality of perspective summarization. Chen et al. (2022b) propose a novel document retrieval paradigm that focuses aggregating and displaying responses from web documents based on varying viewpoints, and reveal that users prefer seeing search results in different clusters of perspectives instead of in a list ranked by relevance. Jiang et al. (2023) tackle opinion summarization for user reviews by generating summaries from different perspectives. Their framework selects subsets of reviews based on sentiment polarity and informational contrast, and produces balanced pros, cons, and verdict summaries. Plepi et al. (2024) demonstrate that a tailored generation model which conditions on a user’s personal context produces more appropriate and empathetic responses than large general models, effectively capturing different viewpoints in the conflict.
To mitigate framing bias and ideology conflicts, Milbauer et al. (2021) propose a method to uncover complex ideological and worldview differences across online communities beyond a single left-vs-right axis. Their study identifies multiple axes of polarization and nuanced ideological distinctions, offering a more multifaceted analysis of online opinion differences. Liu et al. (2022) introduce a pre-training approach for ideology classification that learns ideological bias by directly comparing news articles about the same event reported by outlets with different leanings. Chen et al. (2023) tackle political ideology classification under limited and biased data by disentangling content from style, enabling accurate ideology detection with minimal training data. Lee et al. (2022) present a model that leverages news titles and employs hierarchical multi-task learning to neutralize biased content from title to article, while Liu et al. (2023) induce a neutral event graph that captures events with minimal framing bias by synthesizing across different ideologies.
3 Conflicts in Human-Annotated Texts
3.1 Origins
Annotation Disagreement
The subjective nature of human judgments bring noise and disagreements to their annotated data Kahneman (2021). Such annotation disagreements in NLP arise from multiple sources, including linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices. Uma et al. (2021) provide a comprehensive survey of disagreement across NLP and computer vision tasks, highlighting subjective ambiguity and annotator diversity as key factors of disagreements. Sandri et al. (2023) categorize disagreements in offensive language detection, showing that some stem from inherent ambiguity while others result from annotation errors or lack of context. Their findings imply that not all disagreements are equal – some signal hard-to-classify content, whereas others indicate correctable annotation issues. Jiang and de Marneffe (2022) similarly classify NLI disagreements into three broad categories—linguistic uncertainty, annotator bias, and task design issues. They show that both linguistic uncertainty and annotator bias contribute substantially to label variability, and a significant portion of “disagreement noise” in NLI is systematic and predictable (e.g. stemming from specific ambiguity types or annotator profiles).
Task design also significantly influences annotation disagreements. Dsouza and Kovatchev (2025) find that labels disagreement in Reinforcement Learning from Human Feedback (RLHF) largely depend on annotator selection and task formulation, while Yung and Demberg (2025) show that free-choice annotation methods, where annotators select any suitable connective, yield less diversity (often converging on common labels) than forced-choice approaches, where annotators choose from predefined options, in discourse relation labeling.
On the other hand, demographic and ideological factors also shape disagreement. Pavlick and Kwiatkowski (2019) suggest that many NLI disagreements are not errors but reflect genuine ambiguities in language and individual variations in background knowledge. Sap et al. (2022) demonstrate that annotators’ personal beliefs and identities affect their perception of toxicity, and Wan et al. (2023) find that demographic data significantly enhances the prediction of annotation disagreements.
Ethical and Societal Biases
Ethical and societal biases are also present in human-annotated texts. Biases introduced during annotation, whether related to race, gender, or geography, can significantly skew model predictions and decision-making processes Buolamwini and Gebru (2018). For instance, studies have shown that popular NLP datasets have a severe Western-centric skew Faisal et al. (2022). This Western dominance in training/evaluation data means models optimized on these benchmarks assume a Western context by default. A model might perform well on answering questions about New York or London (since those appear often in the data), but fail on questions about Nairobi or Manila simply due to lack of exposure Faisal et al. (2022).
Sap et al. (2022) show that annotators’ ideological and racial attitudes affect their judgments of toxicity, with conservative annotators less likely to flag anti-Black slurs as toxic but more likely to misclassify African American English (AAE) as offensive. Similarly, Thorn Jakobsen et al. (2022) examine how annotation guidelines interact with annotator demographics, finding that different task formulations can either exacerbate or mitigate biases. They show that even well-designed guidelines can elicit systematically different responses from distinct demographic groups, underscoring the importance of inclusive task framing to reduce disparities in annotations. Pei and Jurgens (2023), on the other hand, introduce POPQUORN, a dataset explicitly designed to measure the impact of annotator demographics across multiple NLP tasks. Their large-scale analysis confirms that annotator background, such as age, gender, race, and education, accounts for substantial variance in annotations.
3.1.1 Implications and Mitigation
Early research has highlighted the impact of annotator disagreements on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al. (2014). Pavlick and Kwiatkowski (2019) find that standard NLI models’ uncertainty did not reflect the true uncertainty found in human opinions. This mismatch suggests that when disagreements are ignored, models become overconfident on contentious cases. In addition, Anand et al. (2023) observe that a classifier provided only with single “gold” labels tends to be less confident and less accurate on examples where annotators widely disagreed on. Such models may treat these instances as “hard to learn” or mislabeled. Furthermore, Sap et al. (2019) reveal how annotator biases can lead to racist outcomes in hate speech classifiers. They uncover a spurious correlation: tweets written in African American English (AAE) are often rated as more toxic by annotators, even when they are not hate speech. Models trained on such data inherit this bias, falsely flagging content by Black authors as offensive at disproportionately high rates.
In terms of data collection, previous work has explored the strategy of acquiring multiple labels for each data item from various annotators to enhance data quality. They develop a probabilistic model to assess the true label of an item by considering the varying expertise of annotators and the possibility of label noise Sheng et al. (2008). Mostafazadeh Davani et al. (2022) examine strategies to train NLP models without discarding annotator disagreements. They propose a multi-annotator modeling approach: a multi-task neural network that learns to predict each individual annotator’s label as a separate output while sharing a common representation. Similarly, different studies have shown that models which incorporate annotator disagreement as "soft" labels (i.e. full label distribution) for training outperform those trained on aggregated single labels Uma et al. (2021); Fornaciari et al. (2021)
4 Conflicts during Model Interactions
4.1 Knowledge Conflicts
4.1.1 Origins
Context vs. Memory
A common type of knowledge conflict arises when a model’s prompt (contextual knowledge) contradicts what the model has learned and stored in its parameters (parametric knowledge) Longpre et al. (2021); Chen et al. (2022a). One prevalent cause of such conflicts is the presence of updated information Chen et al. (2021a); Lazaridou et al. (2021); Luu et al. (2022), where newly available knowledge contradicts models’ previously learned knowledge.
Recent studies have developed many evaluation frameworks and datasets to assess LLMs’ behaviors in this scenario through different methods, including entity substitution Longpre et al. (2021); Chen et al. (2022a); Wang et al. (2024), adversarial perturbation Chen et al. (2022a); Xie et al. (2024), misinformation injection Pan et al. (2023), and machine generation Qian et al. (2023); Ying et al. (2024); Tan et al. (2024).
Within and Across Models
Different models may exhibit conflicts in their knowledge bases. Cohen et al. (2023) investigate how different LLMs can be used to fact-check each other to reveal inconsistencies that imply factually incorrect claims. On the other hand, Zhu et al. (2024) address the issue of inconsistencies between the visual and language components in Large Vision-Language Models (LVLMs). These conflicts arise due to the separate training processes and distinct datasets used for each modality, leading to discrepancies in the knowledge they capture. Even the same model may have internal knowledge conflicts. Zhao et al. (2024) identify intra-model contradictions by paraphrasing the same query multiple times and find answer divergence across different LLMs.
4.1.2 Implications and Mitigation
Interestingly, different studies of knowledge conflicts present seemingly contradictory findings. Some studies claim that models often excessively rely on parametric memory when observing conflicts with contextual knowledge Longpre et al. (2021); Some other studies posit that LLMs tend to ground their answers in retrieved documents in this scenario Chen et al. (2022a); Qian et al. (2023); Tan et al. (2024); or even both – LLMs are highly receptive to context when it is the only evidence presented in a coherent way, but also demonstrate a strong confirmation bias toward parametric memory when both supportive and contradictory evidence to their parametric memory are present Xie et al. (2024).
Various approaches have been proposed to mitigate the consequences of such knowledge conflicts. Longpre et al. (2021) propose a simple method that augments the training set with training examples modified by corpus substitution to mitigate memorization, Chen et al. (2022a) present a calibrator that abstains from predicting on instances with conflicting evidence, and Wang et al. (2024) propose a new instruction-based approach that augment LLMs to first identify knowledge conflicts, then pinpoint conflicting information segments, and lastly provide distinct answers in conflicting scenarios.
4.2 Hallucination
4.2.1 Origins
Factual Hallucinations
Factual hallucinations occur when a model’s output conflicts with real-world facts. TruthfulQA Lin et al. (2022) introduces an adversarial question-answering benchmark, revealing that even the best model at that time (GPT-3) was truthful on only 58% of questions, compared to 94% for humans. Pagnoni et al. (2021) construct the FRANK dataset, which annotates factual errors in summarization, identifying strengths and weaknesses of various metrics. Similarly, Honovich et al. (2021) extend QAGS to knowledge-grounded dialogue using question generation and entailment modeling to assess factual consistency. To further evaluate LLMs’ factual knowledge and reasoning, Hu et al. (2023) introduce Pinocchio, a benchmark comprising 20,713 multiple-choice questions across various domains, timelines, and languages, and assess models on fact composition, temporal reasoning, and adversarial robustness, revealing that LLMs often struggle with factual consistency and are prone to spurious correlations. Mallen et al. (2023) also find that language models struggle with less popular factual knowledge, and that retrieval augmentation helps significantly in these cases.
Contextual Hallucinations
Contextual hallucinations occur when generated text contradicts the given input context, such as in summarization, translation, and generation tasks. Maynez et al. (2020) find that summarization models frequently generate content unfaithful to input documents, with 64% of summaries containing unsupported information. In machine translation, Raunak et al. (2021) analyze hallucinations caused by source perturbations and training noise, and find that slight modifications to input data could trigger off-topic translations. Similarly, Dale et al. (2023) introduce HalOmi, a multilingual benchmark for hallucination and omission detection in machine translation, showing that prior hallucination detectors often fail across different language pairs. In generation tasks, Liu et al. (2021b) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al. (2024) introduce RAGTruth, a comprehensive corpus designed for analyzing word-level hallucinations across various domains and tasks within standard Retrieval-Augmented Generation (RAG) frameworks.
4.2.2 Implications and Mitigation
Mitigating hallucinations in language models has been approached through various strategies, including knowledge disentanglement Neeman et al. (2023), retrieval augmentation Lewis et al. (2020); Shuster et al. (2021), knowledge graphs Guan et al. (2024), and improved verification methods Kryscinski et al. (2020); Wang et al. (2020); Laban et al. (2022); Manakul et al. (2023). DisentQA enhances robustness by training models to separate internal memory from external context, improving accuracy in conflicting knowledge scenarios Neeman et al. (2023). Retrieval-Augmented Generation (RAG) mitigates factual inconsistencies by integrating external sources like Wikipedia Lewis et al. (2020) or incorporating a neural search module into chatbot responses Shuster et al. (2021). In addition, Guan et al. (2024) demonstrate how retrofitting LLM outputs using structured knowledge graphs can correct factual inconsistencies, particularly in complex reasoning tasks. For hallucination detection methods, FactCC and QAGS introduce automated methods using synthetic data and question-answer validation to assess factual consistency Kryscinski et al. (2020); Wang et al. (2020). SummaC refines entailment-based scoring Laban et al. (2022), and SelfCheckGPT detects hallucinations by sampling multiple model outputs and checking for agreement without external references Manakul et al. (2023).
5 Open Challenges
Building Conflict-Aware AI Systems
Conflicts in training data and retrieved contexts can lead to unexpected consequences in NLP models, affecting their reliability and trustworthiness Pan et al. (2023); Liu et al. (2024). Studies show that knowledge conflicts within retrieved contexts severely degrade LLM performance in retrieval-augmented generation (RAG) Chen et al. (2022a); Pham et al. (2024); Liu et al. (2024); Li et al. (2024b), and QA models are highly susceptible to misinformation, where even minimal evidence contamination can lead to incorrect predictions Pan et al. (2023). Also, LLMs exhibit confirmation bias, where models prefer external evidence that aligns with their parametric knowledge, even when conflicting information is present Xie et al. (2024). This bias can cause models to reinforce incorrect or outdated information instead of updating their knowledge based on newly retrieved sources. We highlight the need of developing conflict-aware NLP systems that (1) assess and resolve contradictions in retrieved contexts, (2) mitigate confirmation bias by designing models that reason over conflicting evidence, and (3) integrate fact-checking mechanisms that evaluate source credibility and detect misinformation before generating responses.
Towards Robust and Fair NLP Models
Beyond technical challenges in handling knowledge conflicts, systemic biases in training data and annotation processes impact the fairness of NLP models. The distribution of web data skews LLM outputs toward Western perspectives, disproportionately reflecting Western English-speaking values Ramaswamy et al. (2023); Mihalcea et al. (2024); Tao et al. (2024); Naous et al. (2024). Furthermore, LLMs provide inconsistent answers across languages, revealing contradictions in how they encode cultural perspectives Li et al. (2024a); AlKhamissi et al. (2024). Biases also emerge from annotation processes, where demographic and ideological factors influence how data is labeled Sap et al. (2019). Future directions to create robust and fair NLP models include (1) building demographically diverse and geographically representative training datasets, (2) enhancing consistency of models by aligning cross-lingual knowledge representations, and (3) improving annotation frameworks to account for demographic and ideological biases among annotators.
Limitations
Conflicting information is present both in the data that models rely on and in their generated outputs. While we strive to account for all potential conflict scenarios, some cases may inevitably be overlooked. Additionally, due to space constraints, we do not provide an exhaustive discussion of the literature on each specific type of conflict. Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, patterns, and common challenges.

### v2

@@
 While prior work has addressed some of these conflicts in isolation, we unify them under the broader concept of conflicting information, analyze their implications, and discuss mitigation strategies.
-We highlight key challenges and future directions for developing conflict-aware NLP systems that can reason over and reconcile conflicting information more effectively.
+We highlight key challenges for developing conflict-aware and robust NLP systems, and propose concrete research directions to address them.
 1 Introduction
@@
 (2019).
-Figure 1: Examples of the three different ares of conflicts discussed in this work.
+Figure 1: Examples of the three different areas of conflicts discussed in this work.
 The first example describes a case where two different entities of the same name are found naturally on the web, the second example elaborates the annotation disagreement in a sentiment analysis task, and the third showcases a knowledge conflict between the context and memory of LLMs during model interactions.
-{forest}
-for tree= grow=east, parent anchor=east, child anchor=west, edge=draw, thick, rounded corners, draw=black, fit=band, minimum width=5cm, align=center, anchor=west, s sep=18pt, l sep=22pt, inner sep=10pt [Conflicts in NLP, fill=yellow!40, tier=0 [Model Interactions, fill=yellow!30, tier=1 [Hallucination, fill=green!40, tier=2 [Contextual Hallucination, fill=red!40, tier=3 [Maynez et al.
-(2020),
-Raunak et al. (2021), Dale et al. (2023), fill=cyan!40] ] [Factual Hallucination, fill=red!40, tier=3 [Lin et al. (2022),
-Pagnoni et al. (2021), Honovich et al. (2021), fill=cyan!40] ] ] [Knowledge Conflicts, fill=green!40, tier=2 [Across Different Models, fill=red!40, tier=3 [Cohen et al. (2023),
-Zhu et al. (2024), Zhao et al. (2024), fill=cyan!40] ] [Between Context and Parameters, fill=red!40, tier=3 [Longpre et al. (2021); Chen et al. (2022a),
-Chen et al.
-(2021a); Lazaridou et al.
-(2021), fill=cyan!40] ] ] ] [Human Annotations, fill=yellow!30, tier=1 [Ethical and Societal Biases, fill=green!40, tier=2 [Annotation Biases, fill=red!40, tier=3 [Sap et al.
-(2022), Faisal et al.
-(2022),
-Thorn Jakobsen et al. (2022) , fill=cyan!40] ] ] [Human Judgements, fill=green!40, tier=2 [Interpretation Disagreements, fill=red!40, tier=3 [Kahneman (2021), Uma et al. (2021),
-Sandri et al.
-(2023), Jiang and de Marneffe (2022), fill=cyan!40] ] ] ] [Natural Conflicts on Web, fill=yellow!30, tier=1 [Conflicts in Opinions, fill=green!40, tier=2 [Framing Bias, fill=red!40, tier=3 [Card et al.
-(2015), Liu et al.
-(2019),
-Fan et al. (2019), Lei et al. (2022), fill=cyan!40] ] [Perspectives, fill=red!40, tier=3 [Chen et al. (2019), Liu et al. (2021a),
-Lee et al. (2022) , Wan et al. (2024) , fill=cyan!40] ] ] [Factual Conflicts, fill=green!40, tier=2 [Contradictory Evidence, fill=red!40, tier=3 [Chen et al. (2022a),Hong et al. (2024),
-Liu et al. (2024), Pham et al. (2024), fill=cyan!40] ] [Ambiguity, fill=red!40, tier=3 [Min et al. (2020), Zhang and Choi (2021),
-Dhingra et al. (2022), Cole et al. (2023), fill=cyan!40] ] ] ] ]
-Figure 2: Taxonomy of conflicts in texts.
 Existing work on conflicts in NLP tends to focus on specific issues, such as annotation disagreements Uma et al.
+(2021); Klie et al.
 (2023), hallucinations and factuality Zhang et al.
@@
 In this survey, we conceptualize these diverse challenges under the umbrella of conflicting information and analyze their origins, implications, and mitigation strategies.
-We first examine conflicts inherent in training data, including natural texts from the web and human-annotated datasets.
-We then explore conflicts arising during interactions with models in the LLM era, discussing their impact on downstream tasks and highlighting key challenges and future directions for conflict-aware AI systems.
+{forest}
+t
+Figure 2: Taxonomy of conflicts in texts.
+To ensure comprehensive and representative coverage of conflicts in NLP, we first establish a high-level categorization encompassing three primary sources: (1) natural conflicts present in web data, (2) conflicts arising from human annotation, and (3) conflicts emerging from model interactions.
+Notably, conflicts found in natural web texts and human-annotated datasets are primarily present in the training data—i.e., the inputs to models—whereas conflicts involving model interactions can arise in various forms, such as inconsistencies between model outputs and their inputs, contradictions among multiple outputs, or conflicts within the outputs themselves.
+For each category, we identify influential and widely cited survey papers as initial seed works  Uma et al.
+(2021); Klie et al.
+(2023); Zhang et al.
+(2023); Xu et al.
+(2024); Feng et al.
+(2024); Wang et al.
+(2023).
+Building upon these seeds, we systematically trace and incorporate the most impactful and representative studies for each type of conflict through citation chaining and targeted literature searches across major databases.
+This approach enables us to synthesize developments in each category and connect them, thereby providing an integrated discussion of current challenges, impacts on downstream tasks, and promising future directions for conflict-aware AI systems (See Appendix B for more discussion about our survey methodology).
 The abundance of online data is accompanied by inherent conflicts, stemming from diverse sources, interpretations, and biases.
 These conflicts manifest as factual conflicts, such as semantic ambiguities Pavlick and Tetreault (2016); Min et al.
-(2020) and factual inconsistencies Pham et al.
+(2020) [^1] The previous version on ACL Anthology has a wrong paper title for a reference here.
+and factual inconsistencies Pham et al.
 (2024); Liu et al.
@@
 (2019); Liu et al.
-(2021a).
+(2021).
 Factual conflicts are particularly prevalent in open-domain question answering (QA) and retrieval-augmented generation (RAG) systems Chen et al.
@@
 The presence of conflicting viewpoints complicates tasks such as summarization, sentiment analysis, and dialogue generation, where maintaining coherence and neutrality is crucial Liu et al.
-(2021a); Lee et al.
+(2021); Lee et al.
 (2022).
@@
 (2024).
-By systematically categorizing and analyzing these conflicts, this survey provides a unified perspective on their origins, implications, and mitigation strategies.
-Addressing these issues is essential for building robust and trustworthy AI systems that operate effectively across diverse domains and user groups.
-Our findings contribute to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
+In this survey, we systematically examine the landscape of conflicts in NLP by categorizing them into three primary sources.
+For each conflict type, we detail how such conflicts arise and in what forms they take (origins), the challenges they pose (implications), and the strategies developed to address them (mitigation).
+We present a comprehensive taxonomy in Figure 2, as well as structured summary tables—Table 1, Table 2, and Table 3—that synthesize datasets, methodologies, and analysis from prior work.
+By offering a unified framework for understanding and addressing conflicting information in NLP, this survey contributes to the development of conflict-aware frameworks for data collection, model training, and model usage, ultimately enhancing the fairness and reliability of NLP.
 2 Conflicts in Natural Texts on the Web
@@
 Many other work specifically focus on the temporal aspect of ambiguity, benchmarking and evaluating models’ awareness and adaptation to time-sensitive questions Chen et al.
-(2021b); Liska et al.
+(2021); Liska et al.
 (2022); Kasai et al.
@@
 This issue is particularly prevalent in open-domain question answering settings, where models must navigate inconsistencies across diverse information sources.
-Liu et al.
+For example, Liu et al.
 (2024) find that 25% of unambiguous factual questions queried on Google retrieve conflicting evidence from multiple sources.
-For instance, a Google search for "When was Kendrick Lamar’s first album released?" yields conflicting evidence, illustrating the challenge of integrating contradictory information in retrieval-based QA systems Liu et al.
-(2024).
 Researchers have proposed different datasets to systematically study how NLP models handle such conflicts.
@@
 (2023); Wan et al.
-(2024); Hong et al.
+(2024a); Hong et al.
 (2024), and pre-defined rule-based templates Kazemi et al.
@@
 2.1.2 Implications and Mitigation
-Zhang and Choi (2021) show that pre-trained language models perform competitively at identifying whether a question is context dependent, matching human agreements, but lag behind human-level performance by a significant margin when answering questions dependent on temporal contexts.
-Cole et al.
-(2023) demonstrate that large language models benefit from a “disambiguate-then-answer” pipeline, showing improved reliability by detecting ambiguity before attempting an answer.
-Dhingra et al.
-(2022) propose time-aware language models that condition on timestamps to mitigate confusion arising from outdated facts or evolving knowledge.
-Studies have shown that retrieval performance heavily impacts which sources models rely on Chen et al.
-(2022a), knowledge conflicts within contexts significantly degrade LLMs’ performance in RAG settings Pham et al.
+Implications
+Factual conflicts pose significant challenges for NLP systems.
+Pre-trained language models accurately detect context-dependent questions but fall short when answering queries requiring temporal context, performing notably below human levels Zhang and Choi (2021).
+Additionally, large language models (LLMs) often exhibit confirmation bias, favoring retrieved information that aligns with their parametric memory despite contradictory evidence Xie et al.
+(2024).
+Consequently, conflicting information sources severely impact retrieval-augmented generation (RAG) frameworks, significantly degrading model performance even with minimal misinformation exposure Pham et al.
 (2024); Liu et al.
 (2024); Li et al.
-(2024b), and QA models are vulnerable to even small amounts of evidence contamination brought by misinformation Pan et al.
+(2024b); Pan et al.
 (2023).
-In addition, large language models (LLMs) exhibit a strong confirmation bias, favoring external evidence that aligns with their parametric memory, even when conflicting evidence is present Xie et al.
+Mitigation
+To address these challenges, various mitigation strategies have been proposed.
+Effective methods include fine-tuning calibrators for selective abstention Chen et al.
+(2022a), employing a "disambiguate-then-answer" pipeline to detect ambiguity proactively Cole et al.
+(2023), and developing time-aware models that condition responses on timestamps to manage outdated information Dhingra et al.
 (2022).
-To mitigate these issues, Hong et al.
-(2024) proposed fine-tuning a discriminator or prompting GPT-3.5 to elicit its discriminative capability and show that these approaches significantly enhance model robustness.
-On the other hand, Liu et al.
-(2024) proposed fine-tuning LLMs with human-written explanations to teach models to reason through conflicting evidence.
+Further robustness improvements have been achieved through fine-tuning discriminators or prompting GPT-3.5 models to explicitly recognize conflicting evidence Hong et al.
+(2024), as well as incorporating human-written explanations in fine-tuning processes to enhance models’ reasoning capabilities Liu et al.
+(2024).
 2.2 Conflicts in Opinions
@@
 (2019), posing challenges for downstream tasks like summarization where consolidating viewpoints and presenting unbiased information are crucial Liu et al.
-(2021a); Lee et al.
+(2021); Lee et al.
 (2022).
+Several studies have explored perspectives in the context of conflicting information.
 Chen et al.
-(2019) address this challenge by presenting a range of perspectives on a given claim.
-The authors introduce the task of substantiated perspective discovery, where a system identifies diverse, evidence-supported perspectives that take a stance on a claim, and curated a dataset, PERSPECTRUM, for the task, using online debate platforms and search engines.
+(2019) introduce the task of substantiated perspective discovery, where systems identify diverse, evidence-supported stances on a claim, and release the PERSPECTRUM dataset using online debates and search results.
 Wan et al.
-(2024) introduce ConflictingQA, a dataset comprising controversial questions paired with real-world evidence documents presenting divergent facts, argumentation styles, and conclusions.
+(2024a) propose ConflictingQA, a dataset of controversial questions paired with real-world documents that present divergent facts, arguments, and conclusions.
 Plepi et al.
-(2024) study perspective-taking in contentious online conversations (e.g.
-social media dilemmas).
-The authors create a new corpus of 95k conflict scenarios augmented with each user’s self-disclosed background information.
+(2024) examine perspective-taking in contentious online discourse, curating a corpus of 95k conflict scenarios annotated with users’ self-reported backgrounds.
 Liu et al.
-(2021a) propose MultiOpEd, an open-domain corpus focusing on automatic perspective discovery.
-MultiOpEd comprises 1,397 controversial topics, each accompanied by two editorials expressing opposing viewpoints, and each editorial includes a one-sentence perspective summarizing its core argument and a brief abstract highlighting supporting details.
+(2021) present MultiOpEd, a corpus of 1,397 controversial topics, each paired with opposing editorials and concise summaries capturing their core perspectives.
 Framing Bias
@@
 (2019).
-Various studies have examined different facets of such media bias.
+Numerous studies have investigated different aspects of media bias.
 Card et al.
-(2015) introduce the Media Frames Corpus (MFC) to facilitate the computational study of media framing in news articles.
-The MFC comprises several thousand news articles covering three policy issues, each annotated according to 15 general-purpose framing dimensions.
+(2015) introduce the Media Frames Corpus (MFC), a collection of news articles annotated with 15 general-purpose framing dimensions across three policy issues, enabling computational analysis of media framing.
 Liu et al.
-(2019) introduce Gun Violence Frame Corpus (GVFC), a dataset of news headlines with frames curated and annotated by journalism and communication experts.
+(2019) present the Gun Violence Frame Corpus (GVFC), a dataset of news headlines annotated by domain experts to capture framing in gun violence reporting.
 Fan et al.
-(2019) focus on informational bias, where factual content is subtly framed by content selection and organization within articles.
-They introduce BASIL, a dataset comprising 300 news articles annotated with 1,727 bias spans, revealing that informational bias occurs more frequently than lexical bias.
-Lei et al.
-(2022) identify ideological bias at the sentence level by analyzing news discourse structure, showing that bias-indicative sentences may appear neutral in isolation.
+(2019) examine informational bias—bias conveyed through content selection and structure—and release BASIL, a dataset of 300 news articles annotated with 1,727 bias spans, demonstrating that informational bias is more prevalent than lexical bias.
 2.2.2 Implications and Mitigation
+Implications
 Analysis of PERSPECTRUM reveals significant natural language understanding challenges, as human performance substantially outperforms machine baselines at identifying diverse, evidence-supported perspectives Chen et al.
@@
 Furthermore, when selecting real-world evidence for controversial questions, LLMs predominantly prioritize the relevance of the evidence to the query, often disregarding stylistic attributes such as the presence of scientific references or a neutral tone Wan et al.
-(2024).
+(2024a).
 In addition, the distribution and biases of web data also affects models to behave from a Western perspective Ramaswamy et al.
@@
 (2024a); AlKhamissi et al.
+(2024); Eloundou et al.
 (2025), revealing conflicting cultural perspectives within a single model.
-Several studies have proposed methods to address conflicts in perspectives.
+Mitigation
+Several studies have proposed methods to address conflicts in perspectives and ideological bias.
 Liu et al.
-(2021a) show that incorporating auxiliary tasks enhances the quality of perspective summarization.
+(2021) show that auxiliary tasks improve perspective summarization quality, while Chen et al.
+(2022b) propose a retrieval paradigm that clusters documents by viewpoint, revealing users’ preference for diverse perspectives over relevance-ranked lists.
+Jiang et al.
+(2023b) generate opinion summaries by selecting review subsets based on sentiment polarity and contrast, producing balanced pros, cons, and verdicts.
+Plepi et al.
+(2024) demonstrate that conditioning generation on users’ personal contexts yields more empathetic and appropriate responses than general-purpose models.
+To mitigate framing and ideological bias, Milbauer et al.
+(2021) uncover nuanced worldview differences across communities by identifying multiple axes of polarization beyond the traditional left–right spectrum.
+Liu et al.
+(2022b) pre-train models for ideology detection by comparing reporting on the same events across partisan sources.
 Chen et al.
-(2022b) propose a novel document retrieval paradigm that focuses aggregating and displaying responses from web documents based on varying viewpoints, and reveal that users prefer seeing search results in different clusters of perspectives instead of in a list ranked by relevance.
-Jiang et al.
-(2023) tackle opinion summarization for user reviews by generating summaries from different perspectives.
-Their framework selects subsets of reviews based on sentiment polarity and informational contrast, and produces balanced pros, cons, and verdict summaries.
-Plepi et al.
-(2024) demonstrate that a tailored generation model which conditions on a user’s personal context produces more appropriate and empathetic responses than large general models, effectively capturing different viewpoints in the conflict.
-To mitigate framing bias and ideology conflicts, Milbauer et al.
-(2021) propose a method to uncover complex ideological and worldview differences across online communities beyond a single left-vs-right axis.
-Their study identifies multiple axes of polarization and nuanced ideological distinctions, offering a more multifaceted analysis of online opinion differences.
-Liu et al.
-(2022) introduce a pre-training approach for ideology classification that learns ideological bias by directly comparing news articles about the same event reported by outlets with different leanings.
-Chen et al.
-(2023) tackle political ideology classification under limited and biased data by disentangling content from style, enabling accurate ideology detection with minimal training data.
+(2023) disentangle content from style to enable ideology classification under data scarcity and bias.
 Lee et al.
-(2022) present a model that leverages news titles and employs hierarchical multi-task learning to neutralize biased content from title to article, while Liu et al.
-(2023) induce a neutral event graph that captures events with minimal framing bias by synthesizing across different ideologies.
+(2022) employ hierarchical multi-task learning to neutralize bias from news titles to article bodies, while Liu et al.
+(2023) construct neutral event graphs by synthesizing perspectives across ideological divides.
 3 Conflicts in Human-Annotated Texts
+Conflicts in human-annotated texts largely arise from two sources: annotation disagreements and societal or ethical biases.
+Disagreements stem from linguistic ambiguity, annotator backgrounds, and task design, while biases reflect systematic demographic or ideological influences that can skew labeling in consistent ways.
+Though conceptually distinct, these sources often interact—biases may amplify disagreement or entrench disparities.
+Differentiating between them is essential for understanding annotation-related conflicts and for developing more reliable and equitable NLP datasets.
 3.1 Origins
 Annotation Disagreement
-The subjective nature of human judgments bring noise and disagreements to their annotated data Kahneman (2021).
-Such annotation disagreements in NLP arise from multiple sources, including linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices.
+The subjective nature of human judgment introduces variability and disagreement into annotated data Kahneman (2021).
+In NLP, such disagreements arise from linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices.
 Uma et al.
-(2021) provide a comprehensive survey of disagreement across NLP and computer vision tasks, highlighting subjective ambiguity and annotator diversity as key factors of disagreements.
+(2021) survey disagreements across NLP and vision tasks, identifying subjective ambiguity and annotator diversity as key contributors.
 Sandri et al.
-(2023) categorize disagreements in offensive language detection, showing that some stem from inherent ambiguity while others result from annotation errors or lack of context.
-Their findings imply that not all disagreements are equal – some signal hard-to-classify content, whereas others indicate correctable annotation issues.
-Jiang and de Marneffe (2022) similarly classify NLI disagreements into three broad categories—linguistic uncertainty, annotator bias, and task design issues.
-They show that both linguistic uncertainty and annotator bias contribute substantially to label variability, and a significant portion of “disagreement noise” in NLI is systematic and predictable (e.g.
-stemming from specific ambiguity types or annotator profiles).
-Task design also significantly influences annotation disagreements.
-Dsouza and Kovatchev (2025) find that labels disagreement in Reinforcement Learning from Human Feedback (RLHF) largely depend on annotator selection and task formulation, while Yung and Demberg (2025) show that free-choice annotation methods, where annotators select any suitable connective, yield less diversity (often converging on common labels) than forced-choice approaches, where annotators choose from predefined options, in discourse relation labeling.
-On the other hand, demographic and ideological factors also shape disagreement.
-Pavlick and Kwiatkowski (2019) suggest that many NLI disagreements are not errors but reflect genuine ambiguities in language and individual variations in background knowledge.
+(2023) classify disagreements in offensive language detection as stemming from inherent ambiguity, annotation errors, or contextual gaps, highlighting that some disagreements reflect hard-to-classify content, while others indicate correctable issues.
+Similarly, Jiang and de Marneffe (2022) categorize NLI disagreements into linguistic uncertainty, annotator bias, and task design, showing that much of the observed noise is systematic and predictable.
+Task formulation also plays a critical role.
+Dsouza and Kovatchev (2025) find that label disagreement in reinforcement learning from human feedback (RLHF) is shaped by annotator selection and task phrasing.
+Demographic and ideological factors further influence disagreements.
+Pavlick and Kwiatkowski (2019) argue that many NLI disagreements reflect genuine linguistic ambiguity and individual variation rather than annotation error.
 Sap et al.
-(2022) demonstrate that annotators’ personal beliefs and identities affect their perception of toxicity, and Wan et al.
-(2023) find that demographic data significantly enhances the prediction of annotation disagreements.
+(2022) demonstrate that annotators’ personal beliefs and identities affect toxicity judgments, while Wan et al.
+(2023) show that demographic features significantly improve disagreement prediction.
 Ethical and Societal Biases
-Ethical and societal biases are also present in human-annotated texts.
-Biases introduced during annotation, whether related to race, gender, or geography, can significantly skew model predictions and decision-making processes Buolamwini and Gebru (2018).
-For instance, studies have shown that popular NLP datasets have a severe Western-centric skew Faisal et al.
+Human-annotated texts also encode societal biases related to race, gender, and geography, which can significantly skew model predictions and downstream decisions Buolamwini and Gebru (2018).
+Sap et al.
+(2022) show that annotators’ ideological and racial identities influence toxicity judgments, with conservative annotators less likely to flag anti-Black slurs and more likely to misclassify African American English (AAE) as offensive.
+Thorn Jakobsen et al.
+(2022) examine how annotation guidelines interact with annotator demographics, demonstrating that even well-designed tasks can elicit systematically different responses across groups, highlighting the need for inclusive task design.
+Pei and Jurgens (2023) introduce POPQUORN, a dataset designed to assess demographic effects on annotation across NLP tasks, and find that annotator attributes—such as age, gender, race, and education—account for substantial variance in labeling behavior.
+3.1.1 Implications and Mitigation
+Implications
+Early research has underscored the impact of annotator disagreement on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al.
 (2014).
-This Western dominance in training/evaluation data means models optimized on these benchmarks assume a Western context by default.
-A model might perform well on answering questions about New York or London (since those appear often in the data), but fail on questions about Nairobi or Manila simply due to lack of exposure Faisal et al.
-(2022).
+Pavlick and Kwiatkowski (2019) show that standard NLI models fail to capture the true uncertainty present in human judgments, leading to overconfidence on contentious examples.
+Similarly, Anand et al.
+(2024) find that models trained on single “gold” labels perform poorly and exhibit lower confidence on high-disagreement instances, often treating them as mislabeled or hard to learn.
 Sap et al.
-(2022) show that annotators’ ideological and racial attitudes affect their judgments of toxicity, with conservative annotators less likely to flag anti-Black slurs as toxic but more likely to misclassify African American English (AAE) as offensive.
-Similarly, Thorn Jakobsen et al.
-(2022) examine how annotation guidelines interact with annotator demographics, finding that different task formulations can either exacerbate or mitigate biases.
-They show that even well-designed guidelines can elicit systematically different responses from distinct demographic groups, underscoring the importance of inclusive task framing to reduce disparities in annotations.
-Pei and Jurgens (2023), on the other hand, introduce POPQUORN, a dataset explicitly designed to measure the impact of annotator demographics across multiple NLP tasks.
-Their large-scale analysis confirms that annotator background, such as age, gender, race, and education, accounts for substantial variance in annotations.
-3.1.1 Implications and Mitigation
-Early research has highlighted the impact of annotator disagreements on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al.
-(2014).
-Pavlick and Kwiatkowski (2019) find that standard NLI models’ uncertainty did not reflect the true uncertainty found in human opinions.
-This mismatch suggests that when disagreements are ignored, models become overconfident on contentious cases.
-In addition, Anand et al.
-(2023) observe that a classifier provided only with single “gold” labels tends to be less confident and less accurate on examples where annotators widely disagreed on.
-Such models may treat these instances as “hard to learn” or mislabeled.
-Furthermore, Sap et al.
-(2019) reveal how annotator biases can lead to racist outcomes in hate speech classifiers.
-They uncover a spurious correlation: tweets written in African American English (AAE) are often rated as more toxic by annotators, even when they are not hate speech.
-Models trained on such data inherit this bias, falsely flagging content by Black authors as offensive at disproportionately high rates.
-In terms of data collection, previous work has explored the strategy of acquiring multiple labels for each data item from various annotators to enhance data quality.
-They develop a probabilistic model to assess the true label of an item by considering the varying expertise of annotators and the possibility of label noise Sheng et al.
+(2019) demonstrate how annotator bias can yield discriminatory outcomes: tweets in African American English (AAE) are frequently misclassified as toxic, a bias inherited by models that disproportionately flag content from Black authors.
+Additionally, many widely used NLP datasets exhibit a strong Western-centric skew Faisal et al.
+(2022), causing models to generalize poorly to underrepresented regions—for example, excelling on questions about New York or London, but failing on Nairobi or Manila due to lack of exposure.
+Mitigation
+Prior work has explored collecting multiple labels per data item to capture annotation variability and improve data quality.
+Probabilistic models have been developed to infer true labels by accounting for annotator expertise and label noise Sheng et al.
 (2008).
 Mostafazadeh Davani et al.
-(2022) examine strategies to train NLP models without discarding annotator disagreements.
-They propose a multi-annotator modeling approach: a multi-task neural network that learns to predict each individual annotator’s label as a separate output while sharing a common representation.
-Similarly, different studies have shown that models which incorporate annotator disagreement as "soft" labels (i.e.
-full label distribution) for training outperform those trained on aggregated single labels Uma et al.
+(2022) propose a multi-task neural network that models each annotator’s labels individually while sharing a common representation, preserving disagreement in training.
+Similarly, studies show that models trained on soft labels—i.e., full label distributions reflecting annotator disagreement—consistently outperform those trained on aggregated labels Uma et al.
 (2021); Fornaciari et al.
-(2021)
+(2021).
 4 Conflicts during Model Interactions
+Conflicts during model interactions primarily manifest as knowledge conflicts and hallucinations, each posing distinct challenges.
+Knowledge conflicts occur when a model’s parametric memory contradicts contextual input or when inconsistencies arise across models, whereas hallucinations occur when outputs deviate from real-world facts or the given input.
+Differentiating these two types of conflict clarifies their underlying causes and helps guide targeted mitigation strategies.
 4.1 Knowledge Conflicts
@@
 One prevalent cause of such conflicts is the presence of updated information Chen et al.
-(2021a); Lazaridou et al.
+(2021); Lazaridou et al.
 (2021); Luu et al.
@@
 Within and Across Models
-Different models may exhibit conflicts in their knowledge bases.
+Conflicts may also arise across or within model knowledge bases.
 Cohen et al.
-(2023) investigate how different LLMs can be used to fact-check each other to reveal inconsistencies that imply factually incorrect claims.
-On the other hand, Zhu et al.
-(2024) address the issue of inconsistencies between the visual and language components in Large Vision-Language Models (LVLMs).
-These conflicts arise due to the separate training processes and distinct datasets used for each modality, leading to discrepancies in the knowledge they capture.
-Even the same model may have internal knowledge conflicts.
-Zhao et al.
-(2024) identify intra-model contradictions by paraphrasing the same query multiple times and find answer divergence across different LLMs.
+(2023) explore how different LLMs encode different knowledge and can be used to fact-check one another, uncovering inconsistencies indicative of factual errors.
+Zhu et al.
+(2024) examine cross-modality conflicts in vision-language models, attributing discrepancies between visual and textual components to separate training regimes and distinct data sources.
+Even within a single model, contradictions can emerge: Zhao et al.
+(2024) detect intra-model inconsistencies by paraphrasing queries and observing divergent answers across prompts.
 4.1.2 Implications and Mitigation
+Implications
 Interestingly, different studies of knowledge conflicts present seemingly contradictory findings.
@@
 (2024).
-Various approaches have been proposed to mitigate the consequences of such knowledge conflicts.
+Mitigation
+Several approaches have been proposed to mitigate the impact of knowledge conflicts.
 Longpre et al.
-(2021) propose a simple method that augments the training set with training examples modified by corpus substitution to mitigate memorization, Chen et al.
-(2022a) present a calibrator that abstains from predicting on instances with conflicting evidence, and Wang et al.
-(2024) propose a new instruction-based approach that augment LLMs to first identify knowledge conflicts, then pinpoint conflicting information segments, and lastly provide distinct answers in conflicting scenarios.
+(2021) reduce memorization by augmenting training data through corpus substitution.
+Chen et al.
+(2022a) introduce a calibrator that abstains from prediction when conflicting evidence is detected.
+More recently, Wang et al.
+(2024) propose an instruction-based framework that enables LLMs to identify conflicts, localize conflicting segments, and generate distinct responses for conflicting scenarios.
 4.2 Hallucination
@@
 Factual Hallucinations
-Factual hallucinations occur when a model’s output conflicts with real-world facts.
-TruthfulQA Lin et al.
-(2022) introduces an adversarial question-answering benchmark, revealing that even the best model at that time (GPT-3) was truthful on only 58% of questions, compared to 94% for humans.
+Factual hallucinations arise when a model’s output contradicts real-world facts.
+Lin et al.
+(2022) present TruthfulQA, an adversarial QA benchmark, and show that even top-performing models like GPT-3 were truthful on only 58% of questions, compared to 94% for humans.
 Pagnoni et al.
-(2021) construct the FRANK dataset, which annotates factual errors in summarization, identifying strengths and weaknesses of various metrics.
-Similarly, Honovich et al.
-(2021) extend QAGS to knowledge-grounded dialogue using question generation and entailment modeling to assess factual consistency.
-To further evaluate LLMs’ factual knowledge and reasoning, Hu et al.
-(2023) introduce Pinocchio, a benchmark comprising 20,713 multiple-choice questions across various domains, timelines, and languages, and assess models on fact composition, temporal reasoning, and adversarial robustness, revealing that LLMs often struggle with factual consistency and are prone to spurious correlations.
+(2021) construct FRANK, a dataset for identifying factual errors in summarization, while Honovich et al.
+(2021) extend QAGS to dialogue by leveraging question generation and entailment for factual consistency evaluation.
+To assess factual knowledge and reasoning in LLMs, Hu et al.
+(2024) introduce Pinocchio, a large benchmark covering multiple domains, timelines, and languages, revealing challenges in composition, temporal reasoning, and robustness.
 Mallen et al.
-(2023) also find that language models struggle with less popular factual knowledge, and that retrieval augmentation helps significantly in these cases.
+(2023) further find that models struggle with less common factual knowledge, with retrieval augmentation significantly improving performance in such cases.
 Contextual Hallucinations
@@
 In generation tasks, Liu et al.
-(2021b) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al.
+(2022a) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al.
 (2024) introduce RAGTruth, a comprehensive corpus designed for analyzing word-level hallucinations across various domains and tasks within standard Retrieval-Augmented Generation (RAG) frameworks.
 4.2.2 Implications and Mitigation
+Implications
+Hallucinations threaten trust, safety, and the integrity of AI-powered workflows.
+Hallucinated outputs can rapidly spread false information.
+For instance, in 2023, an AI-generated image purporting to show an explosion near the Pentagon went viral, briefly causing public panic and even a stock market dip before being debunked Sun et al.
+(2024).
+Hallucinations directly degrade the performance of downstream applications like abstractive summarization.
+Studies have found that a large portion of generated summaries contain unsupported facts, misleading readers and propagating misinformation in news and scientific dissemination Kryscinski et al.
+(2020).
+Mitigation
 Mitigating hallucinations in language models has been approached through various strategies, including knowledge disentanglement Neeman et al.
@@
 (2023).
-5 Open Challenges
-Building Conflict-Aware AI Systems
-Conflicts in training data and retrieved contexts can lead to unexpected consequences in NLP models, affecting their reliability and trustworthiness Pan et al.
-(2023); Liu et al.
-(2024).
-Studies show that knowledge conflicts within retrieved contexts severely degrade LLM performance in retrieval-augmented generation (RAG) Chen et al.
-(2022a); Pham et al.
-(2024); Liu et al.
-(2024); Li et al.
-(2024b), and QA models are highly susceptible to misinformation, where even minimal evidence contamination can lead to incorrect predictions Pan et al.
-(2023).
-Also, LLMs exhibit confirmation bias, where models prefer external evidence that aligns with their parametric knowledge, even when conflicting information is present Xie et al.
-(2024).
-This bias can cause models to reinforce incorrect or outdated information instead of updating their knowledge based on newly retrieved sources.
-We highlight the need of developing conflict-aware NLP systems that (1) assess and resolve contradictions in retrieved contexts, (2) mitigate confirmation bias by designing models that reason over conflicting evidence, and (3) integrate fact-checking mechanisms that evaluate source credibility and detect misinformation before generating responses.
-Towards Robust and Fair NLP Models
-Beyond technical challenges in handling knowledge conflicts, systemic biases in training data and annotation processes impact the fairness of NLP models.
-The distribution of web data skews LLM outputs toward Western perspectives, disproportionately reflecting Western English-speaking values Ramaswamy et al.
+5 Connections, Challenges and Directions
+Given the significance and impact of conflicts in NLP, we advocate for increased attention to the development of conflict-aware and robust AI systems.
+In this section, we highlight specific challenges by connecting different types of conflicts and propose concrete research directions to address them.
+Culturally Robust LLMs
+Among the challenges outlined in this survey, the development of culturally robust LLMs remains particularly underexplored.
+Cultural conflicts emerge both in naturally occurring web data and human-annotated datasets, where Western-centric distributions dominate.
+Prior studies reveal that LLMs often reflect the values and perspectives of Western, English-speaking populations Ramaswamy et al.
 (2023); Mihalcea et al.
@@
 (2024); Naous et al.
+(2024), with misalignments especially pronounced for underrepresented personas and culturally sensitive topics Al Kuwatly et al.
 (2020).
-Furthermore, LLMs provide inconsistent answers across languages, revealing contradictions in how they encode cultural perspectives Li et al.
+Additionally, LLMs exhibit inconsistent behavior across languages Li et al.
 (2024a); AlKhamissi et al.
-(2024).
-Biases also emerge from annotation processes, where demographic and ideological factors influence how data is labeled Sap et al.
-(2019).
-Future directions to create robust and fair NLP models include (1) building demographically diverse and geographically representative training datasets, (2) enhancing consistency of models by aligning cross-lingual knowledge representations, and (3) improving annotation frameworks to account for demographic and ideological biases among annotators.
+(2024); Eloundou et al.
+(2025), revealing internal cultural conflicts.
+These issues are rooted in the data: both the pre-train data and benchmark datasets commonly exhibit Western-centric biases Mihalcea et al.
+(2025); Faisal et al.
+(2022), causing models to default to Western contexts and perform poorly on less-represented regions and cultures.
+However, to the best of our knowledge, no effective methodology has yet been proposed to address this issue.
+With the emergence of culturally distinct LLMs—such as Qwen, trained largely on Chinese data Bai et al.
+(2023), and Vikhr, trained on Russian data Nikolich et al.
+(2025)—a promising direction is model fusion across culturally diverse models to achieve greater cultural balance Wan et al.
+(2024b); Jiang et al.
+(2023a).
+Furthermore, advances in culture-specific LLMs and synthetic data generation offer the potential to curate more culturally representative training and evaluation datasets beyond Western-centric narratives, supporting the development of culturally robust LLMs.
+Building Conflict-Aware AI Systems
+As outlined in this survey, various types of conflicts can arise in a model’s input, each requiring different handling depending on the task.
+We argue that downstream applications should not treat all conflicts uniformly; rather, responses should be tailored to the conflict type.
+For instance, conflicts due to ambiguity should elicit clarification questions, factual contradictions should trigger reasoning over evidence, and opinion-based disagreements should induce balanced, multi-perspective responses.
+Realizing such capabilities requires models to be aware of the potential conflicts and classify them according to a systematic taxonomy.
+Yet, current research lacks frameworks to distinguish and operationalize these conflict types.
+Our proposed taxonomy offers a foundational step toward enabling conflict-aware systems that can recognize, interpret, and appropriately address diverse conflicts in downstream applications.
+6 Conclusion
+We present a unified view of conflicting information in NLP, organizing the landscape into conflicts originating from (i) natural texts on the web, (ii) human annotations, and (iii) model interactions.
+This taxonomy connects lines of work that are often studied in isolation and clarifies how conflicts arise, what they imply for reliability, and how current methods aim to mitigate them.
+Our synthesis argues that conflict awareness should guide the full pipeline: data collection that preserves disagreement and multiple perspectives, models that detect and categorize conflicts and respond with clarification, reasoning, or balanced presentation, and evaluations that measure calibration under disagreement, robustness to contradictory evidence, and cultural coverage across languages and regions.
 Limitations
@@
 While we strive to account for all potential conflict scenarios, some cases may inevitably be overlooked.
-Additionally, due to space constraints, we do not provide an exhaustive discussion of the literature on each specific type of conflict.
-Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, patterns, and common challenges.
+Additionally, due to space constraints, we cannot provide an exhaustive discussion of the literature on each specific type of conflict.
+Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, challenges, and future directions.
+Acknowledgment
+This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via 2022-22072200003.
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
+This work is also supported by NSF grant #IIS-2135581.
+Appendix A Summary Tables
+In the summary tables, dataset covers prior work that proposed datasets and benchmark, method covers work that focus on mitigation strategies, and analysis presents work that aim at providing insights through experiments.
+Table 1: Datasets, methods, and analysis for conflicts in natural texts
+Conflict Type Sub-type Category Work
+Factual Ambiguity Dataset SituatedQA Zhang and Choi (2021)
+AmbigQA Min et al. (2020)
+Time-sensitive QA Chen et al. (2021)
+StreamingQA Liska et al. (2022)
+Real-time QA Kasai et al. (2023)
+Method Disambiguate then answer Cole et al. (2023)
+Time-aware LM Dhingra et al. (2022)
+Contradictory Evidence Dataset QACC Liu et al. (2024)
+Contra-Doc Li et al. (2024b)
+WhoQA Pham et al. (2024)
+Machine-generated Pan et al. (2023)
+Machine-generated Wan et al. (2024a)
+Machine-generated Hong et al. (2024)
+Machine-generated Jiayang et al. (2024)
+Entity-substitution Chen et al. (2022a)
+Rule-based Kazemi et al. (2023)
+Method Finetuned Calibrator Chen et al. (2022a)
+Finetuned w/ Explanation Liu et al. (2024)
+Finetuned discriminator Hong et al. (2024)
+Analysis Confirmation bias Xie et al. (2024)
+Opinion Perspectives Dataset PERSPECTRUM Chen et al. (2019)
+Multi-OpEd Liu et al. (2021)
+NeuS Lee et al. (2022)
+ConflictingQA Wan et al. (2024a)
+Reddit Plepi et al. (2024)
+Method Multi-task learning Liu et al. (2021)
+Opinion summarization Jiang et al. (2023b)
+Tailored generation Plepi et al. (2024)
+Framing Bias Dataset MFC Card et al. (2015)
+GVFC Liu et al. (2019)
+BASIL Fan et al. (2019)
+Method Multifaceted analysis Milbauer et al. (2021)
+Pre-training Liu et al. (2022b)
+Disentanglement Chen et al. (2023)
+Multi-task learning Lee et al. (2022)
+Analysis Sentence-level Lei et al. (2022)
+Table 2: Datasets, methods, and analysis for conflicts in human-annotated texts
+Conflict Type Sub-type Category Work
+Human-Annotated Disagreement Dataset Twitter Sandri et al. (2023)
+RLHF Dsouza and Kovatchev (2025)
+DiscoGeM Yung and Demberg (2025)
+NLI Pavlick and Kwiatkowski (2019)
+Method Probabilistic model Sheng et al. (2008)
+Multi-task Mostafazadeh Davani et al. (2022)
+Soft labels Uma et al. (2021)
+Soft labels Fornaciari et al. (2021)
+Analysis Survey Uma et al. (2021)
+Survey Klie et al. (2023)
+Offensive language Sandri et al. (2023)
+NLI Jiang and de Marneffe (2022)
+Task design Dsouza and Kovatchev (2025)
+Free choice Yung and Demberg (2025)
+Personal belief Sap et al. (2022)
+Demographic data Wan et al. (2023)
+Biases Dataset Gender Buolamwini and Gebru (2018)
+Argument mining Thorn Jakobsen et al. (2022)
+POPQUORN Pei and Jurgens (2023)
+Analysis Western-centric Faisal et al. (2022)
+Toxicity Wan et al. (2023)
+Racist outcome Sap et al. (2019)
+Table 3: Datasets, methods, and analysis for conflicts during model interactions
+Conflict Type Sub-type Category Work
+Knowledge Context vs. Memory Dataset Entity substitution Longpre et al. (2021)
+Entity substitution Chen et al. (2022a)
+Instruction-based Wang et al. (2024)
+Misinformation injection Pan et al. (2023)
+KRE Ying et al. (2024)
+context-conflicting Tan et al. (2024)
+Method Data Augmentation Longpre et al. (2021)
+Abstention Chen et al. (2022a)
+Instruction-based Wang et al. (2024)
+Within & Across Analysis LM-vs-LM fact-checking Cohen et al. (2023)
+Cross-modality Zhu et al. (2024)
+Intra-model contradiction Zhao et al. (2024)
+Hallucination Factual Dataset TruthfulQA Lin et al. (2022)
+FRANK Pagnoni et al. (2021)
+$q^{2}$ Honovich et al. (2021)
+Pinocchio Hu et al. (2024)
+MiniCheck Tang et al. (2024)
+Method RAG Lewis et al. (2020)
+RAG Shuster et al. (2021)
+Knowledge graph Guan et al. (2024)
+Disentanglement Neeman et al. (2023)
+QA validation Kryscinski et al. (2020)
+QA validation Wang et al. (2020)
+Entailment-based Laban et al. (2022)
+SelfCheckGPT Manakul et al. (2023)
+Analysis Less popular entities Mallen et al. (2023)
+Contextual Dataset HalOmi Dale et al. (2023)
+HADES Liu et al. (2022a)
+RAGTruth Niu et al. (2024)
+Method Context-aware decoding Shi et al. (2024)
+Long context Liu et al. (2025)
+Context-DPO Bi et al. (2024)
+CR-DPO Huang et al. (2025)
+Analysis Summarization Maynez et al. (2020)
+Translation Raunak et al. (2021)
+Appendix B Survey Methodology
+Our objective is to map the broad and diverse landscape of textual conflicts in NLP, identify central implications and gaps, and outline a research agenda.
+We surveyed research indexed in Google Scholar, the ACL Anthology, and OpenReview, covering peer-reviewed work in major ML and NLP venues such as ICLR and ACL, as well as recent manuscripts on arXiv.
+Rather than relying on predetermined keyword filters, we primarily used citation chaining: starting from influential surveys corresponding to the three sources of conflict considered in this work (natural web text, human annotations, and model interactions), we applied both backward and forward chaining using Google Scholar and venue indexes.
+Paper selection did not follow a rigid checklist.
+The two authors independently screened and included papers based on inclusion criteria that prioritized recency, citation impact, and perceived influence on the area, then reconciled disagreements through discussion.
+This methodology emphasizes coverage and structure over exhaustiveness and is intended to synthesize a rapidly evolving field while making the scope of inclusion explicit and surfacing open problems.